What is the answer of ordered pairs for
y=x+3
y=2x+5

Answer by jim_thompson5910(35256)

Start with y = x+3 and plug in y = 2x+5. Then solve for x

y = x+3
2x+5 = x+3
2x = x+3-5
2x-x = 3-5
x = -2

------------
Now that we know x = -2, use this to find y.

y = x+3
y = -2+3
y = 1

-------------------------

Since we got x = -2 and y = 1, the final answer is (-2,1)